A skier slides down from the top of a frictionless 400 incline of height 205 m. If she starts from rest, calculate the speed of the skier at the bottom of the incline.
Answer #1

Solution:

Given:

Height of the incline (h) = 205 m

Since the inclined plane is friction-less.

According to conservation of energy: KEfinal = PEinitial

(1/2) mv2 = m g h

v2 = 2 g h

v2 = (2)(9.81 m/s2)(205 m) = 4022.1

v = 63.42 m/s

Thus: the speed of the skier at the bottom of the incline is : v = 63.42 m/s